# Text-to-SQL Task-oriented Dialogue Ontology Construction

### v1

@@ 4.2 Evaluation
We run TeQoDO on a dataset, generating the ontology by executing all update queries per dialogue. For evaluation, tables map to domains, columns to slots, and values to ground truth values. System actions and user intents are matched by aligning table names with the “system actions” “and user intents” keys and comparing values directly to the ground truth.
We adapt the evaluation framework of Lo et al. (2024) to our task-oriented dialogue ontology structure, which includes domains, slots, values, system actions, and user intents. Specifically, we use literal F1 as a hard metric and fuzzy and continuous F1 as soft metrics, as these capture most relevant evaluation aspects Lo et al. (2024). The hard metric captures syntactic similarity, respectively, while the soft metrics capture some human intuition on semantic similarity. Since these metrics treat all ontology graph edges equally — overlooking infrequent higher-level relations — we extend them to account for hierarchical levels and their specific edge types.
@@
In the literal score, only exact term matches count: true positives are $V_{i_{\textup{pred}}}\cap V_{i_{\textup{true}}}$, false positives are $V_{i_{\textup{pred}}}\setminus V_{i_{\textup{true}}}$, and false negatives are $V_{i_{\textup{true}}}\setminus V_{i_{\textup{pred}}}$.
@@
In both the fuzzy and continuous setups, we map nodes above a similarity threshold to ground truth nodes. We use a sentence transformer and set the threshold at $t_{\textup{{sim}}}=0.436$ to consider two concepts equivalent as in (Lo et al., 2024). For the fuzzy setup, all predicted nodes above the threshold are mapped to ground truth. True positives are predicted nodes with cosine similarity above the threshold to at least one ground truth node, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\}$. False positives are predicted nodes with similarity below or equal to the threshold for all ground truth nodes, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\forall v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$. False negatives are ground truth nodes without any predicted node mapped above the threshold, i.e., $\{v_{t}\in V_{i_{\textup{true}}}\mid\forall v_{p}\in V_{i_{\textup{pred}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$.
@@
For the continuous metric, only the predicted node with the highest similarity to each ground truth node above $t_{\textup{sim}}$ is a true positive, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\,|\,\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\land\operatorname{sim}(v_{p},v_{t})=\max_{v_{p}^{\prime}\in V_{i_{\textup{pred}}}}\operatorname{sim}(v_{p}^{\prime},v_{t})\}.$ This stricter metric penalises multiple predictions for the same ground truth node. We use this as the primary metric, as it accounts for surface-form variations without allowing significantly different structures — an approach that, through qualitative analysis, proved to reward the best ontologies in terms of downstream usability.
We match the hierarchy top-down: domains first, then slots, values, intents, and actions. Only slots of matched domains are considered; unmatched domains exclude their slots. This applies similarly to values, intents, and actions.
On Wikipedia and ArXiv, comparison results are from Lo et al. (2024) using their evaluation scripts. We report their literal, fuzzy, continuous, and graph F1 metrics. Graph F1 embeds predicted and ground truth graphs, comparing their structure through graph convolutions.
4.3 Models
DORE (Vukovic et al., 2024)
stands for dialogue ontology relation extraction, which uses Gemma-2B instruct (Team et al., 2024) to predict dialogue-level ontology relations between known terms in the prompt. DORE enhances transfer learning via constrained chain-of-thought decoding (Wang and Zhou, 2024), restricting decoding to known terms and relations. Final outputs are chosen based on confidence across sampled responses. Table˜1 presents the best-performing DORE models on both TOD datasets, including fine-tuned and transfer models, trained on the respective other dataset.
GenDSI (Finch et al., 2024)
is the generative dialogue state inference approach (GenDSI). It fine-tunes T5-3B (Raffel et al., 2020) on a large synthetic TOD dataset (Finch and Choi, 2024) to generate dialogue state updates from user-system turns. State updates, as slot-values, are clustered, with the most frequent slot name in each cluster chosen as representative. For evaluation, domains are extracted from slots using the first word of each slot name to form the ontology hierarchy. Predictions lacking domain names are discarded. As DORE and GenDSI do not predict user intents or system actions, evaluation is limited to domains, slots, and values, with averages calculated accordingly in Table˜1.
@@ 4.3 Models
Table 1: Ontology construction comparison. DORE (Vukovic et al., 2024) and GenDSI (Finch et al., 2024) do not predict the intents and actions; hence, their average is only over domains, slots, and values.
OLLM (Lo et al., 2024)
is fine-tuned on Wikipedia and ArXiv ontologies and uses a custom masking loss and rule-based aggregation of article-level predictions. On these datasets, we also compare to Hearst patterns Roller et al. (2018), which use hand-crafted lexico-syntactic rules, e.g. “mammals such as humans”, to predict is-a relations. REBEL Huguet Cabot and Navigli (2021) treats relation extraction as a translation task and trains a language model to generate relations.
TeQoDO
We use GPT-4o-mini (“gpt-4o-mini-2024-07-18”) OpenAI (2024) for all experiments. The model performs well on HumanEval Du et al. (2024), which includes SQL coding tasks. Full pipeline prompts appear in Section˜A.1. SQLite [^3] https://www.sqlite.org is used via Python’s sqlite3.444https://docs.python.org/3/library/sqlite3.html Final prompts were manually crafted according to best practices [^5] https://help.openai.com/en/articles/6654000-best-practices-for-prompt-
engineering-with-the-openai-api
and refined with ChatGPT (see Section˜B.2 for prompt variant performance).
For more consistency, we sample values for each table column to expose the LLM to concrete instances in the column value example set-up.
To reduce variation in column values (e.g. “Alexander Bed and Breakfast” vs “Alexander B & B”) and improve naming consistency in the DB, we also experiment with query results for similar tables, columns, and values. We call this similarity matching. This decreases duplicate entries by increasing the chance of reusing existing values. The model tends to use the LIKE operator in SQLite, which matches substrings but fails in cases like the example above. Instead, we compute semantic similarity using the “all-MiniLM-L6-v2” sentence-transformers model Reimers and Gurevych (2019) and a similarity threshold of 0.4360.4360.436. This threshold, found by Lo et al. (2024), corresponds to the median similarity for synonyms in WordNet Miller (1994) using the same model. For each query, we return up to 5 similar concepts that exceed this threshold.
@@ 4.4 Task-oriented Ontology Construction
Table˜1 shows that DORE overfits to domains and slots in the training set but outperforms TeQoDO on value-level and literal metrics. This stems from supervised fine-tuning, enabling DORE to learn exact value phrasing. GenDSI surpasses DORE on domains and slots but is outperformed by TeQoDO on all metrics. GenDSI’s lack of explicit domain prediction causes slot predictions to be dropped if domains are not the first word in slot names. These results show that TeQoDO predicts higher-level hierarchical concepts far better than the other methods.
Ablation
Table˜2 shows the ablation study results, while Section˜B.1 details results for domains, slots, values, intents, and actions.
In the table, direct update refers to generating updates without querying the database. Query update involves querying the database before applying updates. DST Step indicates the use of the DST step. Success indicates that user goal fulfilment is included in the update prompt (Section˜3.2). Sim. denotes the use of similarity matching. Ex. refers to column value examples (Section˜4.3).
The large discrepancy between literal and similarity-based metrics aligns with the findings of Lo et al. (2024), highlighting numerous surface form variations that literal evaluation fails to capture. For example, a domain named “hotel_bookings” is a false positive in literal evaluation despite being a reasonable prediction, since the ground truth domain is “hotel”. Comparing fuzzy and continuous metrics, precision on the latter is lower on both datasets, as only one of several predictions can match each ground truth concept.
Using the query and update pipeline improves all scores on both datasets, already surpassing DORE in the more expressive similarity-based metrics Lo et al. (2024). This is expected, as allowing the model to query existing tables leads to better alignment and more consistent update queries. Without querying first, the model tends to create new table names for every dialogue, reducing precision. Recall is generally higher in the direct update approach due to generating more tables, but this substantially lowers precision.
@@
Table 2: TOD ontology construction ablation study for MultiWOZ and SGD test sets. We report the macro average scores and standard deviation for 5 seeds of different dialogue orders over the five hierarchy classes: domains, slots, values, system actions, and user intents. Bold F1 scores are significantly better than Query Update ($p<0.05$). On SGD Query Update approaches, we input batches of 10 dialogues for faster inference.
@@
Table 4: Zero-shot DST results for TripPy-R (Heck et al., 2022) with ground truth and TeQoDO induced ontology inference on MultiWOZ 2.1 in joint goal accuracy per domain.
On both datasets, incorporating dialogue theory improves performance and reduces variance. The DST step alone does not significantly outperform the query and update baseline. On MultiWOZ, similarity matching, column value examples, and success yield significantly better performance. On SGD, using success together with column value examples significantly improves the continuous metric. Note that the combination of similarity matching and success yields no improvements, and is therefore omitted for brevity. Dialogue theory narrows the gap between fuzzy and continuous metrics, indicating more consistent DB updates. The influence of dialogue order is notably diminished, as evidenced by the drop in variance when dialogue theory is applied.
@@
Incorporating dialogue success into the DB update prompt significantly reduces erroneous SQL query ratios, achieving single-digit error rates on both datasets. This suggests that applying dialogue theory improves the model’s handling of the database and generation of SQL update queries, as most errors arise from incorrect column names, indicating poor schema handling.
@@ 4.5 Downstream Application: DST
Approach Literal F1 Fuzzy F1 Continuous F1 Graph F1 Supervised
Wikipedia
TeQoDO 0.03 $\dagger$64.94 34.76 $\dagger$64.15
Hearst Patterns (Roller et al., 2018) 0.30 53.80 35.00 54.40
REBEL (Huguet Cabot and Navigli, 2021) $\dagger$0.40 62.40 $\dagger$35.60 7.20
OLLM (Lo et al., 2024) 9.30 91.50 50.00 64.40
ArXiv
TeQoDO + Clustering 0.00 $\dagger$33.95 $\dagger$34.15 89.89
Hearst Patterns (Roller et al., 2018) 0.00 0.00 15.10 55.30
REBEL (Huguet Cabot and Navigli, 2021) 0.00 6.00 28.10 54.60
OLLM (Lo et al., 2024) 4.00 57.00 35.70 $\dagger$63.30
Table 5: Ontology construction results for Wikipedia and ArXiv test sets from Lo et al. (2024). We use the metrics from their code base and the Hearst, REBEL and OLLM results from their work. In bold the best F1 score for each column is highlighted and with a $\dagger$ the second highest.
To demonstrate downstream usability, we apply the TeQoDO-induced ontology in a specialised dialogue state tracking model. We replicate the zero-shot leave-one-domain-out setup from Heck et al. (2022). Using their TripPy-R model, we train with one domain held out and infer on that domain using the TeQoDO-induced ontology instead of ground truth. Only domain and slot predictions are used in this setup.
@@
4.6 General Ontology Construction
When applying TeQoDO to general ontologies, we truncate the prompt from the left to fit the model’s context window, needed only for Wikipedia. Wikipedia’s table count reaches thousands, indicating that pre-filtering tables in TeQoDO’s first step may be necessary—this is left for future work.
On ArXiv, we evaluate only the first two hierarchy levels, as the ground truth ontology contains just two. Due to the dataset’s size and the small target ontology, directly applying TeQoDO results in too many predicted tables. To address this, we cluster table names and select representative tables based on their proximity to each centroid. This clustering approach mirrors OLLM’s frequency-based aggregation but avoids hand-crafted rules. We use $k$-means MacQueen et al. (1967), choosing $k$ via the silhouette score Rousseeuw (1987), with $k$ ranging from 5 to 20. To speed up inference, we prompt with batches of 5 articles for ArXiv and 200 for Wikipedia.
Results
See Table˜5 for the final TeQoDO results compared to models from Lo et al. (2024) (see Section˜A.4 for prediction and ground truth examples). On Wikipedia, TeQoDO does not outperform OLLM on most metrics, likely due to OLLM’s extensive supervised training. However, TeQoDO surpasses REBEL in fuzzy and graph F1, where it performs comparably to OLLM. On the continuous metric, TeQoDO performs on par with Hearst patterns and REBEL.
By clustering and merging tables, we reduce predicted table numbers and achieve competitive results on ArXiv. TeQoDO attains the highest graph F1 and second-best continuous F1 on ArXiv, outperforming fine-tuned and few-shot models and matching supervised OLLM. This shows TeQoDO’s induced ontologies closely match the ground truth structure. In fuzzy and continuous F1, TeQoDO surpasses Hearst patterns and REBEL. Lower literal F1 scores reflect the absence of supervision in TeQoDO. Results suggest text-to-SQL can adapt to ontologies beyond TOD, though ArXiv requires adjustments. The differing hierarchy depth poses challenges, as SQL suits the three-level TOD structure best. We argue that ArXiv’s high-level labels, with only 61 nodes for thousands of abstracts, underrepresent its content richness.
@@ Results
Table 6: Qualitative comparison between the query and update baseline and proposed improvements. Note that the hotel tables were named differently in the different runs of the ablations.
The main challenges for TeQoDO on general ontologies are the dataset size and differing hierarchy levels compared to TOD ontologies. However, the graph F1 results indicate that the structure can be effectively induced on much larger ontologies, showing the generalisability of our approach.
4.7 TOD Ontology Construction Qualitative Analysis
When prompting the model to generate an SQL DB, the induced hierarchy aligns more closely with the desired structure than when explicitly prompting for domains, slots, and values. The latter leads to more misclassifications, such as predicting slots as domains.
See Table˜6 for a qualitative comparison of update queries on MultiWOZ test dialogues, detailed in Section˜A.3. In the similarity matching example, only “Allenbell” appears in the dialogue, while “The Allenbell” is stored from a previous dialogue. Without similarity matching, the model queries “Allenbell”, yielding no result due to the missing “The”. Similarity matching retrieves “The Allenbell” from the database. The model chooses which results to include, as less related concepts may also be retrieved.

### v2

@@ 4.2 Evaluation
We run TeQoDO on a dataset, generating the ontology by executing all update queries per dialogue. For evaluation, tables map to domains, columns to slots, and values to ground truth values. System actions and user intents are matched by aligning table names with the “system actions” and “user intents” keys that are explicitly stored in the ground truth. This means that the model is expected to predict one table for system actions and one for user intents. The values are then directly compared to the values stored under the keys in the ground truth ontology.
We adapt the general ontology evaluation framework of lo2024ontologylearning to our task-oriented dialogue ontology structure, which includes domains, slots, values, system actions, and user intents. Specifically, we use literal F1 as a hard metric and fuzzy and continuous F1 from lo2024ontologylearning as soft metrics, as these capture most relevant evaluation aspects according to their results. The hard metric captures syntactic similarity, while the soft metrics capture some human intuition on semantic similarity. Since these metrics treat all ontology graph edges equally – overlooking infrequent higher-level relations – we extend them to account for hierarchical levels and their specific edge types by evaluating the different levels separately. We use all metrics for the ablation study and choose the continuous metric as our main metric, as it covers semantic similarity while being strict on the final structure of the ontology. Note that we adopt the naming of the chosen metrics from lo2024ontologylearning to remain aligned with established terminology from general ontology construction to facilitate a comparison.
@@
Here, only exact matches count: true positives are $V_{i_{\textup{pred}}}\cap V_{i_{\textup{true}}}$, false positives are $V_{i_{\textup{pred}}}\setminus V_{i_{\textup{true}}}$, and false negatives are $V_{i_{\textup{true}}}\setminus V_{i_{\textup{pred}}}$.
@@
In both the fuzzy and continuous setups, we map nodes above a cosine similarity threshold to ground truth nodes. We use a sentence transformer reimers-gurevych-2019-sentence and set the threshold at $t_{\textup{{sim}}}=0.436$. [^4] This threshold corresponds to the median similarity for synonyms in WordNet miller-1994-wordnet; lo2024ontologylearning using the “all-MiniLM-L6-v2” sentence-transformers model. For the fuzzy setup, all predicted nodes above the threshold are mapped to ground truth. True positives are predicted nodes that are mapped to at least one ground truth node, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\}$. False positives are predicted nodes with similarity below or equal to the threshold for all ground truth nodes, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\mid\forall v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$. False negatives are ground truth nodes without any predicted node mapped above the threshold, i.e., $\{v_{t}\in V_{i_{\textup{true}}}\mid\forall v_{p}\in V_{i_{\textup{pred}}}:\operatorname{sim}(v_{p},v_{t})\leq t_{\textup{sim}}\}$.
@@
For the continuous metric, only the predicted node with the highest similarity to each ground truth node above $t_{\textup{sim}}$ is a true positive, i.e., $\{v_{p}\in V_{i_{\textup{pred}}}\,|\,\exists v_{t}\in V_{i_{\textup{true}}}:\operatorname{sim}(v_{p},v_{t})>t_{\textup{sim}}\land\operatorname{sim}(v_{p},v_{t})=\max_{v_{p}^{\prime}\in V_{i_{\textup{pred}}}}\operatorname{sim}(v_{p}^{\prime},v_{t})\}.$ This stricter metric penalises multiple predictions for one ground truth node. We use this as the main metric, as it accounts for surface-form variations without allowing significantly different structures – an approach that, through qualitative analysis, proved to reward the best ontologies in terms of downstream usability. We match the hierarchy top-down: domains first, then slots, values, intents, and actions. Only slots of matched domains and values of matched slots are considered.
Graph F1
On Wikipedia and arXiv, comparison results are from lo2024ontologylearning using their evaluation scripts. We report their literal, fuzzy, continuous, and graph F1 metrics. Graph F1 measures structural similarity between a predicted graph $G^{\prime}$ and a gold graph $G$. Each node is embedded using a 2-layer graph convolution, and nodes are matched by cosine similarity. Let $s_{\text{graph}}$ be the total similarity of the best node matching. Graph precision and recall are defined as: $\text{Graph-Precision}=\frac{s_{\text{graph}}}{|V^{\prime}|},\text{Graph-Recall}=\frac{s_{\text{graph}}}{|V|},$ where $V^{\prime}$ and $V$ are the node sets of $G^{\prime}$ and $G$, respectively. For more details, see lo2024ontologylearning.
4.3 Baselines
As DORE and GenDSI do not predict user intents or system actions, evaluation is limited to domains, slots, and values in Table˜1.
DORE (vukovic-etal-2024-dialogue)
is defined as dialogue ontology relation extraction and fine-tunes Gemma-2B Instruct (team2024gemma) to predict dialogue-level ontology relations between terms in the prompt. Transfer learning is enhanced through constrained chain-of-thought decoding (wang2024chain), which selects the final response according to confidence across sampled responses while restricting decoding to known terms. Table˜1 shows the best-performing DORE models, including fine-tuned models and transfer models.
@@ 4.3 Baselines
Table 1: Ontology construction comparison with continuous metric F1. DORE (vukovic-etal-2024-dialogue) and GenDSI (finch-etal-2024-transforming) do not predict the intents and actions; hence, their average is only over domains, slots, and values. We re-run both baselines using their published code bases.
GenDSI (finch-etal-2024-transforming)
is the generative dialogue state inference approach. T5-3B (raffel2020t5) is fine-tuned on a large synthetic TOD dataset (finch-choi-2024-diverse) to generate dialogue-state updates from user–system turns. Slot–value updates are clustered, and the most frequent slot name in each cluster is selected as the representative. Domains are extracted from slot names by taking their first word to construct the ontology hierarchy for evaluation and comparison. Predictions lacking domain names are discarded.
OLLM (lo2024ontologylearning)
is a Mistral 7B model jiang2023mistral7b fine-tuned on Wikipedia and arXiv ontologies using a custom masking loss to make article-level ontology relation predictions. These predictions are aggregated with frequency-based rules to form the final ontology.
Hearst patterns roller-etal-2018-hearst
use hand-crafted lexico-syntactic rules, e.g. “mammals such as humans”, to predict is-a relations on Wikipedia and arXiv. Additionally, lo2024ontologylearning apply the smoothed relation matrix to weigh relations among the gold concepts.
REBEL huguet-cabot-navigli-2021-rebel-relation
treats relation extraction as translation and trains a language model to generate relations on Wikipedia and arXiv. lo2024ontologylearning keep only relations of type “subclass of”, “instance of”, “member of”, and “part of” and apply low-rank smoothing.
4.4 TeQoDO
We use GPT-4o-mini (“gpt-4o-mini-2024-07-18”) gpt4o_blog_openai_2024 for all experiments. The model performs well on HumanEval du2021humaneval, which includes SQL coding tasks. Full pipeline prompts are in Section˜A.1. SQLite (Footnote˜2) is used via Python’s sqlite3.555python sqlite3 Final prompts were manually crafted according to best practices [^6] OpenAI Prompting Best Practices and refined with ChatGPT (Section˜4.10).
In the column value example set-up, we sample values for each table column to expose the LLM to concrete instances for more consistency. In the similarity matching set-up, we compute semantic similarity using the “all-MiniLM-L6-v2” sentence-transformers model reimers-gurevych-2019-sentence and a similarity threshold of $0.436$ (Footnote˜4). For each query, we return up to 5 similar concepts that exceed this threshold, which is more versatile than using the LIKE operator from SQL for instance.
For general ontology data, we update the prompt to account for differing hierarchy levels. We prompt the model to generate a DB capturing categories with parent IDs to represent more than three hierarchy levels. A possible SQL representation uses columns for the category name and parent relation. This follows the prompt from lo2024ontologylearning. We aggregate the ontology graph using category names and parent IDs to build hierarchical relations. This ontology graph can be used directly in the lo2024ontologylearning evaluation script. On the arXiv dataset, we cluster top-level categories to account for high-level labels.
@@ 4.5 Task-oriented Ontology Construction
In Table˜1 we see the continuous F1 evaluation of TeQoDO compared to baselines. It shows that DORE overfits to domains and slots in the training set but outperforms TeQoDO on value-level and literal metrics. This stems from supervised fine-tuning, enabling DORE to learn exact value phrasing. GenDSI surpasses DORE on domains and slots but is outperformed by TeQoDO on all metrics. GenDSI’s lack of explicit domain prediction causes slot predictions to be dropped if domains are not the first word in slot names. These results show that TeQoDO predicts higher-level hierarchical concepts far better than the other methods.
@@
Table 2: Ablation study for MultiWOZ and SGD test sets with macro averages and standard deviations for 5 random dialogue orders over the five hierarchy classes. Bold F1 scores are significantly better than Query Update ($p<0.05$). On SGD, we input batches of 10 dialogues to Query Update approaches.
Ablation
@@
Table˜2 shows the ablation study results, while Section˜B.1 details results for domains, slots, values, intents, and actions. In the table, direct update means generating a database update directly without first querying the existing database. Query update means generating an update only after querying the DB. DST Step indicates that the approach includes a form of DST, where the model predicts the structured state before forming the update. Success indicates that the prompt explicitly includes the user’s goal or success condition. Sim. denotes similarity matching, where DB query results are expanded based on semantic similarity to increase the chance of results regardless of phrasing. Ex. refers to column value examples, where the column-information query result includes column–value pairs to guide update generation (See Sections˜3.2 and 4.4 for details).
The discrepancy between literal and similarity-based metrics aligns with the findings of lo2024ontologylearning, literal evaluation fails to capture surface form variations. For example, a domain named “hotel_bookings” is a false positive in literal evaluation despite being a reasonable prediction for the domain “hotel”. Precision on the continuous metric is lower than fuzzy, because only one prediction can match each ground truth concept.
@@
Table 4: Zero-shot DST results for TripPy-R (heck-etal-2022-robust) using ground truth vs TeQoDO ontology on MultiWOZ 2.1 in JGA per domain.
Using the query and update pipeline improves all scores on both datasets, already surpassing DORE in the more similarity-based metrics that are semantically more expressive according to lo2024ontologylearning. This is expected, as allowing the model to query existing tables leads to better alignment and more consistent update queries. Without querying first, the model creates new table names for every dialogue, reducing precision. Recall is higher in the direct update approach due to more tables, but this substantially lowers precision.
Approach Literal F1 Fuzzy F1 Continuous F1 Graph F1 Unsupervised
Wikipedia
TeQoDO 0.03 64.94 34.76 64.15
Hearst Patterns (roller-etal-2018-hearst) 0.30 53.80 35.00 54.40
REBEL (huguet-cabot-navigli-2021-rebel-relation) 0.40 62.40 35.60 7.20
OLLM (lo2024ontologylearning) 9.30 91.50 50.00 64.40
arXiv
TeQoDO + Clustering 0.00 33.95 34.15 89.89
Hearst Patterns (roller-etal-2018-hearst) 0.00 0.00 15.10 55.30
REBEL (huguet-cabot-navigli-2021-rebel-relation) 0.00 6.00 28.10 54.60
OLLM (lo2024ontologylearning) 4.00 57.00 35.70 63.30
Table 5: Ontology construction results for Wikipedia and arXiv test sets from lo2024ontologylearning. We use the metrics from their code base and the Hearst, REBEL, and OLLM results from their work. In bold the best F1 score for each column is highlighted and the second highest in italics.
On both datasets, incorporating concepts from modular TOD systems improves performance and reduces variance. The DST step alone does not significantly outperform the query and update baseline. On MultiWOZ, similarity matching, column value examples, and success yield significantly better performance. On SGD, using success together with column value examples significantly improves the continuous metric. Note that the combination of similarity matching and success yields no improvements, and is therefore omitted for brevity. The concepts from modular TOD systems narrow the gap between fuzzy and continuous metrics, indicating more consistent DB updates. Furthermore, the influence of dialogue order is diminished, as seen by the drop in variance.
@@
Incorporating dialogue success into the DB update prompt significantly reduces erroneous SQL query ratios, achieving single-digit error rates on both datasets. This suggests that applying concepts from modular TOD systems improves the model’s handling of the DB and SQL update queries, as most errors arise from incorrect column names, indicating poor schema handling.
@@ 4.6 Downstream Application: DST
To show downstream usability, we apply the TeQoDO-induced ontology in a specialised dialogue state tracking model. We replicate the zero-shot leave-one-domain-out setup from heck-etal-2022-robust. We train the TripPy-R model with one domain held out on the human-made ground truth ontology and do inference on that domain using the TeQoDO-induced ontology instead of ground truth. We use joint goal accuracy (JGA) as metric.
The mapping uses continuous evaluation, ensuring that at most one domain–slot pair is assigned to each ground truth slot. During DST training, human-labelled slot names are replaced with predicted slots that match them. If no predicted slot reaches the similarity threshold for a ground truth slot, that slot is excluded from training, and the model can never predict it, so its DST output remains empty. For DST evaluation, predicted slots are mapped back to the ground truth slots.
@@
Table 6: Qualitative comparison between the query and update baseline and proposed improvements.
4.7 General Ontology Construction
For general ontologies, we left-truncate the prompt to fit the model’s context window. This can remove earlier instructions and intermediate results from the final-step context. Because Wikipedia contains thousands of tables, pre-filtering in TeQoDO’s first step may be necessary, which we leave for future work.
On arXiv, we evaluate only the first two hierarchy levels, as the ground truth contains only two. Given the dataset’s size and the small target ontology, directly applying TeQoDO yields too many predicted tables. We therefore cluster table names and select representative tables by their proximity to each centroid. Unlike OLLM’s frequency-based aggregation, this removes the need for hand-crafted rules. We use $k$-means macqueen1967some and choose $k$ via the silhouette score ROUSSEEUW198753, with $k$ ranging from 5 to 20. To speed up inference, we input batches of $5$ articles for arXiv and $200$ for Wikipedia.
The challenges for TeQoDO on general ontologies are dataset size and differing hierarchy depths compared to TOD ontologies. Hence, such ontologies are stored in SQL with category names and parent IDs for deeper hierarchies. These tables translate directly into an ontology graph by linking categories according to their parent IDs. The resulting graph is evaluated against the ground-truth using lo2024ontologylearning code.
Results
See Table˜5 for TeQoDO compared to models from lo2024ontologylearning (see Section˜A.4 for examples). On Wikipedia, TeQoDO does not outperform OLLM on most metrics, likely due to OLLM’s extensive supervised training. However, TeQoDO surpasses REBEL in fuzzy and graph F1, where it performs comparably to OLLM. On the continuous metric, TeQoDO performs on par with Hearst patterns and REBEL.
Approach LLaMA 4 Scout (109B/17B) Qwen3 (235B/22B) Llama 3.1 (8B) GPT-OSS (20B/3.6B) GPT-4o-mini
Spider multi-turn text-to-SQL exact match
Text-to-SQL 27.1 29.0 25.9 22.4 40.2
MultiWOZ Continuous F1
Direct Updates $21.4_{\pm 5.5}$ $32.6_{\pm 14.5}$ $22.8_{\pm 6.0}$ $21.9_{\pm 3.0}$ $19.2_{\pm 1.0}$
DB Query Update $28.8_{\pm 11.9}$ $32.8_{\pm 6.3}$ $22.5_{\pm 3.0}$ $43.8_{\pm 6.2}$ $50.1_{\pm 6.7}$
+ DST, Ex., Succ. $33.4_{\pm 5.4}$ $51.9_{\pm 7.1}$ $25.3_{\pm 3.2}$ $40.2_{\pm 5.6}$ $65.2_{\pm 4.7}$
SGD Continuous F1
Direct Updates $39.6_{\pm 4.9}$ $30.6_{\pm 1.1}$ $18.5_{\pm 1.3}$ $47.8_{\pm 11.6}$ $18.6_{\pm 2.3}$
DB Query Update $54.3_{\pm 7.9}$ $60.3_{\pm 7.3}$ $41.4_{\pm 8.3}$ $68.3_{\pm 5.0}$ $44.6_{\pm 21.1}$
+ DST, Ex., Succ. $63.2_{\pm 4.0}$ $67.7_{\pm 3.9}$ $55.7_{\pm 4.1}$ $57.8_{\pm 4.9}$ $61.6_{\pm 5.4}$
Table 7: Continuous F1 scores and standard deviations of open models of different sizes and GPT-4o-mini with 5 random dialogue-order seeds. For MoE models, we indicate both the total parameter count and the active parameters per forward pass. On SGD, we use batches of 10 dialogues to speed up inference. Additionally, we show the multi-turn exact match performance on the Spider dataset yu-etal-2018-spider.
By clustering tables, obtain competitive results on arXiv. TeQoDO achieves the highest graph F1 and second-best continuous F1 on arXiv, outperforming fine-tuned and few-shot models and matching supervised OLLM. This indicates that TeQoDO’s induced ontologies closely match the ground truth structure. In fuzzy and continuous F1, TeQoDO surpasses Hearst patterns and REBEL. Lower literal F1 scores follow from the absence of supervision in TeQoDO. It adapts to ontologies beyond TOD, though arXiv requires specific adjustments. The differing hierarchy depth poses challenges, as SQL fits the three-level TOD structure best. We argue that arXiv’s high-level ontology, with only 61 nodes for thousands of abstracts, underrepresents its content richness.
4.8 TOD Ontology Qualitative Analysis